Verify that each equation is an identity.
2cos^2(theta)-1= 1-tan^2(theta)/1+tan^2(theta)

Answer by lwsshak3(11628)

Verify that each equation is an identity.
2cos^2(theta)-1= 1-tan^2(theta)/1+tan^2(theta)
**
Start with right side:
1-tan^2(theta)/1+tan^2(theta)
[1-(sin^2/cos^2)]/[1+(sin^2/cos^2)]
[(cos^2-sin^2)/cos^2]/[(cos^2+sin^2)/cos^2]
cos^2 cancels out and cos^2+sin^2=1
cos^2-sin^2
=cos^2-(1-cos^2)
=cos^2-1+cos^2)
=2cos^2-1
verified: right side=left side